OpenStudy (anonymous):

Write a compound inequality to represent all of the numbers between -4 and 6. Write your answer using proper set notation.

OpenStudy (anonymous):

are -4 and 6 also included

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I think it's -4 < x < 6 but I'm not 100% sure.

OpenStudy (anonymous):

Acc. to your question \[-4 \le x \le 6\] representing in set notation \[x \epsilon \left[ -4,6 \right]\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

np!